Update for the Week Ending April 1, 2016
Site Remediation Continues
- Continued the application of water and calcium chloride for dust control;
- Continued the application of odor-suppressant foam and BioSolve for dust and odor control;
- Maintained erosion and sediment (E&S) controls (Installed waddles - wood chip and straw filled fabric tubes - around the Site perimeter to prevent silt and sediment from leaving the Site and covered all stockpiles);
- Continued operation of the water treatment plant (WTP) and discharge to the sewer;
- Continued the full scale In-Situ Stabilization (ISS) Bucket Soil Mixing (BSM) in the Holder No. 2 tank;
- Continued pre-drilling for deep obstruction identification and removal along the alignment of the support of excavation (SOE) wall;
- Continued crushing reusable concrete and soil for backfill in the deep excavation; and
- Continued the lining, loading, and shipment of trucks of impacted soils to Bayshore Soil Management for treatment and disposal:
- 14 trucks; 346.67 tons shipped last week; and
- 1,768.99 tons shipped to date.
Recently completed activities:
- Completed the ISS BSM stabilization of five cells in the Holder No. 2 tank. Twenty of twenty cells have been stabilized to date.
- Complete the ISS in the Holder No. 2 tank with the stabilization, above the twenty stabilized cells, of the remaining, top material ;
- Continue the lining, loading, and shipment of trucks of impacted soil to Bayshore Soil Management for treatment and disposal;
- Ship removed asphalt for recycling; and
- Complete pre-clearing and initiate the installation of the cement/bentonite slurry trench for installation of the SOE wall.
Community Air Monitoring Plan Summary
Community air monitoring, initiated on Tuesday, December 1, 2015, continues as intrusive (ground disturbance) activities are conducted.
There were no fifteen-minute Action Limit exceedances Total Volatile Organic Compounds (TVOCs) or odors during this reporting period.
There were three instances of the fifteen-minute average concentration of Particulate Matter (PM-10) [dust] exceeding the applicable Action Limit. Dust was measured above the Action Limit of 150 µg/m3 at AMS-2 on March 28, 2016 caused by spraying water to clean the grout plant near AMS-2. Dust was measured at the Action Limit of 150 µg/m3 at AMS-3 on March 31, 2016 and April 1, 2016 due to a rock crusher blowing small amounts of dust in the air. Additional water was applied with a fire hose during crushing, and the dust levels decreased below the Action Limit.
If you have any questions or comments about the work on the Site, please use the Williamsburg MGP Telephone Hotline at (718) 403-3053.
Photos of Activity
Crushing clean rock and reuse material and applying water for dust control.
Pre-drilling for deep obstruction identification along the SOE wall alignment.
Conducting ISS BSM in the Holder No. 2 tank beneath the temporary containment building (TCB).
Applying long duration foam to the reuse soil stockpiles for dust control.